OpenStudy (anonymous):

Let(-8,3) be a point on the terminal side of theta . Find the exact values of sin theta ,csc theta , and cot theta sin theta = csc theta = cot theta =

OpenStudy (anonymous):

So sin theta = 3/sqrt(73) csc theta = sqrt(73)/3 cot theta = -8/3 is my answer correct ?

OpenStudy (debbieg):

They all look right to me! Just keep in mind for the first one, your teacher might want you to rationalize that den'r. (I would. :) But yes, your method for computing them is correct. :)